# Measuring Computational Universality of Fully Homomorphic Encryption

### v1

@@ 6.4. Private Decision Tree Evaluation
Private Decision Tree Evaluation (PDTE) (Akavia et al., 2022; Kiss et al., 2019; Akhavan Mahdavi et al., 2023; Azogagh et al., 2022) allows a server to provide predictions using a private decision tree on a client’s confidential input, preserving both input privacy (the server learns nothing about the client’s data) and model privacy (the client only obtains the inference result without learning any details of the decision tree).
Decision trees primarily involve comparisons and traversals (Song and Ying, 2015). In these trees, comparisons are made between the client’s inputs and decision thresholds, yielding ciphertext values $c_{i}\in[0,1]$ at each node, while traversals determine the active path by computing the product $c_{1}\cdot c_{2}\cdot\dots\cdot c_{d}$, where $c_{i}$ is the comparison result at the $i$-th level and $d$ is the tree’s depth. The correct decision is identified by the leaf node that decrypts to 1, with all other leaves decrypting to 0. Since private decision tree evaluation involves both linear and non-linear operations, it requires universal FHE methods. Specifically, linear operations following non-linear comparisons correspond to workload-2 defined in Section 5. Notably, the traversal operation $c_{1}\cdot c_{2}\cdot\dots\cdot c_{d}$ can be reformulated (Akhavan Mahdavi et al., 2023) as:
@@
Experimental Analysis. In Figure 5 (a), we evaluate the performance of various universal FHE methods on a complete decision tree with a depth of 6 across different input bit lengths. Figure 5 (b) shows the performance for complete decision trees of varying depths using an 8-bit input. The results indicate that for deeper trees, the word-wise Encoding Switching method performs more efficiently than others since it enables SIMD so that can compare the same input with multiple thresholds simultaneously while avoiding the significant cost of scheme switching between bit-wise and word-wise ciphertexts. On the other hand, the word-wise Scheme Switching perform worst because the significant frequent of switching since every non-linear operation following a linear operation.
@@ 6.5. Private Sorting
Private sorting is a protocol that enables secure ordering of encrypted data without revealing the underlying values or the sorting process. In this scenario, a client encrypts its data and sends to a server, which then performs the sorting directly on the ciphertexts. Private sorting is critical and commonly used in privacy sensitive area such as finacial. For instance, several financial institutions might want to merge and sort transaction data to detect market trends without exposing individual client details.
The most efficient homomorphic sorting algorithm in terms of runing time is the direct sorting algorithm proposed by Çetin, et al. (Çetin et al., 2015). This approach directly determines final position of each element by counting the number of elements less than it. Specifically, for an array of size $m$, the server performs $m(m-1)/2$ homomorphic less-than operations and an equal number of homomorphic additions on the less-than operations’ result for computing the Hamming weight, which yields the correct indices $I$ for each element.
@@
where $S[i]$ is the element at the position $I$ of sorted array $S$ and $X$ is the input array. $\text{EQ}(i,I[j])$ is 1 if the $i$ equals the index $I[i]$ corresponding to $X[j]$, and 0 otherwise. Constructing one position $S[i]$ requires $m$ equality evaluation and $m$ multiplication, so sorting the entire encrypted array requires $O(m^{2})$ non-linear operations (less-than and equality evaluation) and $O(m^{2})$ linear operations (addtions and multiplications).
Thus, homomorphically sorting an encrypted array of $m$ elements demands universal FHE methods, particularly as the linear operations are performed on the results of non-linear operations, corresponding to our workload-2.
@@
Experimental Analysis. We conduct experiments using various universal FHE methods on the private sorting tasks. Figure 7 (a) presents the performance of different FHE methods under different input bit precisions, and Figure 7 (b) shows the performance under different array dimensions. For instance, sorting an 8-element array requires 54.4 seconds with TFHE, compared to 645 seconds for Encoding Switching and 7,062 seconds for Scheme Switching. This efficiency advantage is due to two key factors. First, unlike other applications, private direct sorting on a single array cannot leverage SIMD (Iliashenko and Zucca, 2021), which diminishes the benefits of word-wise methods like Scheme Switching and Encoding Switching. Second, in this context, one of the operands in the linear operations is the output of a non-linear operation (either 0 or 1), resulting in a simpler circuit compared to linear operations on arbitrary integers. Conversely, Scheme Switching shows the lowest efficiency because its frequent switching, triggered by each non-linear operation following a linear one, adds significant overhead.
@@
The reasons are two-fold; firstly, unlike other applications, private direct sorting an encrypted array cannot leverage the SIMD (Iliashenko and Zucca, 2021), which disables the advantages of word-wise Scheme Switching and Encoding Switching. Secondly, one of the linear operation operands is the result of non-linear operation, i.e., 0 or 1. So that the TFHE circuit would be much less than linear operation between two random nature integers. On the other hand, Scheme Switching presents the lowest efficiency across all input bits and array sizes because of the frequent switching since every non-linear operation will follow a linear operation.
@@ 6.6. Private Database Aggregation
Private Database Aggregation (PDBA) is a protocol that protects client privacy during aggregate queries on cloud-stored databases, such as Microsoft Azure SQL Server (mir, 2022a) and AWS Aurora (AWS, 2022). Recently, Leidos has partnered with AWS to explore the solution of using FHE to protect database (AWS, 2023). It allows clients to securely store encrypted data on a cloud server and perform operations, like summation, without disclosing data or the details of the queries. The server processes these functions on encrypted data and returns only the encrypted aggregate results, ensuring the server does not access actual data or query specifics.
Recently, researchers have developed encrypted databases to ensure data privacy while enabling encrypted data processing. Cryptography-based solutions (Popa et al., 2011; Tu et al., 2013; Papadimitriou et al., 2016; Hackenjos et al., 2020), utilize cryptographic primitives such as DET (Bellare et al., 2007), SE (Cash et al., 2013; Curtmola et al., 2006), partially PHE (Paillier, 1999; Rivest et al., 1978) and order OPE/ORE (Chenette et al., 2016; Boldyreva et al., 2009). Recent advancements have explored the application of FHE in HEDA (Ren et al., 2022), HE3DB (Bian et al., 2023), ArcEDB (Zhang et al., 2024a) and Engorgio (Bian et al., 2025).
@@
In this query, the filtering predicates "salary * work_hours BETWEEN 5000 AND 6000" and "salary + bonus BETWEEN 700 AND 800" involve a non-linear operation following a linear operation. This requirement necessitates universal FHE methods and aligns with workload-1 defined in Section 5.
While word-wise FHE methods work well for linear operations, they are less effective for non-linear operations like comparisons and logic filtering. To address this limitation and leverage the high efficiency of SIMD, Ren et al. introduce a novel framework HEDA (Ren et al., 2022). The core idea is to utilize scheme switching between RLWE and LWE ciphertexts. This approach conducts non-linear operations on LWE ciphertexts, where such operations are more naturally supported. Then, it converts the results back to RLWE ciphertexts to perform linear operations in a SIMD fashion.
Experimental Analysis. As described in Section 6.6, SQL queries involve combinations of non-linear and linear operations. To evaluate the performance of each universal FHE method on the private database query, we conducted experiments on Query 1 across databases with different numbers of rows under 8-bit precision, as illustrated in Figure 8 (b).
The results indicate that the word-wise universal FHE methods generally outperform the bit-wise TFHE, largely due to their use of SIMD capabilities, a feature that the bit-wise TFHE lacks. Furthermore, as the number of rows increases, Encoding Switching demonstrates better efficiency than Scheme Switching. For example, Encoding Switching takes 115 seconds for a 512-row database, compared to 119 seconds for Scheme Switching. This advantage is attributed to Encoding Switching’s ability to perform non-linear operations in a SIMD-style, while Scheme Switching processes non-linear operations on bit-wise ciphertexts and cannot leverage SIMD.
@@ 6.7. Private Neural Network
One of the most valuable applications of FHE is in facilitating privacy-preserving deep learning, i.e., Deep Learning as a Service (DLaaS) (Srinivasan et al., 2019; Mohassel and Zhang, 2017), which is commonly used in information sensitive areas such as finacial (Han et al., 2019) and healthcare (Kim and Lauter, 2015). This scenario involves a server that possesses a deep learning model, i.e., neural network, and a client who encrypts their input data before sharing with the untrusted server. The server processes encrypted input to provide predictions without accessing any sensitive information, thereby ensuring a secure inference service.
@@
Neural networks consist of linear layers such as convolutions and non-linear layers like ReLU and MaxPool (Krizhevsky et al., 2012; LeCun et al., 1998). The combination of linear and non-linear operations required by these layers necessitates universal FHE methods; indeed, the iterative sequence of linear and non-linear layers forms the basis of workload-3 as defined in Section 5.
@@
Despite the use of universal FHE methods, some research leveraged the specific feature of neural network to improve efficiency but sacrificing some accuracy. Specifically, polynomial approximation-based methods (Lee et al., 2022a, 2023b, b) leverage the noise resilience of neural networks to enhance efficiency while reducing multiplicative depth and complexity. Although these approximations are typically accurate only within a narrow range (often between -1 and 1), activation layer inputs are usually constrained (e.g., within $[-32,32]$ or $[-64,64]$). By scaling these inputs to fit within $[-1,1]$ using a predetermined scalar, servers can effectively implement these approximation techniques throughout the network, as demonstrated in approaches like DaCapo (Cheon et al., [n. d.]), OPT-CNN (Kim and Guyot, 2023), and Hetal (Lee et al., 2023a) among others (Lee et al., 2022c; Dathathri et al., 2020). However, determining the appropriate scaling factor is critical: if it is too large, inputs may not be sufficiently constrained within $[-1,1]$; if too small, the inputs become overly diminished, resulting in significant approximation errors near 0.
@@
Another approximation strategy simulates non-linear activation functions using linear operations. For example, CryptoNets (Gilad-Bachrach et al., 2016) replaces ReLU with the square function, a method later adopted by frameworks like LoLa (Brutzkus et al., 2019) and Falcon (Lou et al., 2020a). This approach avoids the overhead of homomorphic non-linear operations by training the neural network with a substitute activation function.
@@
Experimental Analysis. To evaluate the performance of various private neural network inference implementations described in Section 6.7, we conducted experiments using the CIFAR-10 dataset (Krizhevsky et al., 2009) on a VGG-9 model. VGG-9 is a streamlined version of the VGG network family, comprising 9 layers including convolutional and fully connected layers. Typically, it consists of 6 convolutional layers followed by 3 fully connected layers, with ReLU activations and max pooling operations interleaved between the convolutional layers. Encryption parameters were set as specified in each referenced paper to ensure at least 128-bit security.
@@
Figure 8 (a) shows that implementations using the square function as an activation substitute, such as Falcon (Lou et al., 2020a) and LoLa (Brutzkus et al., 2019), yield lower performance. This is primarily because training with degree-2 polynomial activations can lead to instability issues like exploding or vanishing gradients (Ali et al., 2020; Xue et al., 2022). Although the square function enables faster inference, it is generally only suitable for smaller, less complex neural networks.
@@
In contrast, private neural networks that employ approximation-based activation functions, such as OPP-CNN (Kim and Guyot, 2023) and EVA (Dathathri et al., 2020), achieve higher accuracy levels. Meanwhile, TFHE’s performance is significantly slower compared to implementations using interpolation approximations. This discrepancy arises because neural network architectures predominantly consist of linear layers, which benefit from the SIMD capabilities of word-wise CKKS schemes. Bit-wise TFHE, lacking SIMD support and optimized linear operations, exhibits lower performance. Additionally, the Scheme Switching method (SHE (Lou and Jiang, 2019)) achieves the highest accuracy but at slower inference times than the approximation-based approaches.
@@ 7. Recommendation and Discussion
First, if the application involves only linear operations, word-wise FHE methods (BGV, BFV, CKKS) are the most efficient choice. Conversely, if the application requires only non-linear operations, the next consideration should be whether the application can benefit from the word-wise SIMD technique. If SIMD does not offer an advantage, a bit-wise FHE method, such as TFHE, is the most efficient option. However, if SIMD is useful, several techniques enable non-linear functionalities in a word-wise context, including direct polynomial interpolation (Narumanchi et al., 2017), decomposition-based polynomial interpolation (Iliashenko and Zucca, 2021), general functional bootstrapping (Geelen and Vercauteren, 2023), or special encoding methods like XCMP (Akhavan Mahdavi et al., 2023). These word-wise methods enable parallel computation to improve efficiency.
When an application demands both linear and non-linear computations (i.e., universality-required applications), the next question is whether the workload can benefit from SIMD parallelism. If SIMD does not offer advantages, e.g., private sorting for an array, a bit-wise FHE method, such as TFHE, remains the most practical choice. If SIMD is beneficial, the final consideration is whether exact, error-free results are necessary. If approximate results are acceptable, as is common in tasks like neural network inference, where noise tolerance is inherent, CKKS with approximation (Cheon et al., 2020; Lee et al., 2021) offers the most efficiency. If exact computation is required, the Encoding Switching (Zhang et al., 2023) between beFV and FV, and the Scheme Switching (Lu et al., 2021) between word-wise and bit-wise FHE can deliver the best performance.

### v2

@@ 7.4. Private Decision Tree Evaluation
Private Decision Tree Evaluation (PDTE) (akavia2022privacy; kiss2019sok; akhavan2023level; azogagh2022probonite) allows a server to provide predictions using a private decision tree on a client’s confidential input, preserving both input privacy (the server learns nothing about the client’s data) and model privacy (the client only obtains the inference result without learning any details of the decision tree).
Decision trees primarily involve comparisons and traversals (song2015decision). In these trees, comparisons are made between the client’s inputs and decision thresholds, yielding ciphertext values $c_{i}\in[0,1]$ at each node, while traversals determine the active path by computing the product $c_{1}\cdot c_{2}\cdot\dots\cdot c_{d}$, where $c_{i}$ is the comparison result at the $i$-th level and $d$ is the tree’s depth. The correct decision is identified by the leaf node that decrypts to 1, with all other leaves decrypting to 0. Since private decision tree evaluation involves both linear and non-linear operations, it requires general FHE methods. Specifically, linear operations following non-linear comparisons correspond to workload-2 defined in Section 6. Notably, the traversal operation $c_{1}\cdot c_{2}\cdot\dots\cdot c_{d}$ can be reformulated (akhavan2023level) as:
@@
Experimental Analysis. In Figure 5 (a), we evaluate the performance of various general FHE methods on a complete decision tree with a depth of 6 across different input bit lengths. Figure 5 (b) shows the performance for complete decision trees of varying depths using an 8-bit input. The results indicate that for deeper trees, the word-wise Encoding Switching method performs more efficiently than others since it enables SIMD so that it can compare the same input with multiple thresholds simultaneously while avoiding the high cost of scheme switching between bit-wise and word-wise ciphertexts. On the other hand, the word-wise Scheme Switching performs worst because of the significant frequency of switching, since every non-linear operation follows a linear operation.
@@ 7.5. Private Sorting
Private sorting is a protocol that enables secure ordering of encrypted data without revealing the underlying values or the sorting process. In this scenario, a client encrypts its data and sends it to a server, which then performs the sorting directly on the ciphertexts. Private sorting is critical and commonly used in privacy-sensitive areas such as financial. For instance, several financial institutions might want to merge and sort transaction data to detect market trends without exposing individual client details.
The most efficient homomorphic sorting algorithm in terms of running time is the direct sorting algorithm proposed by Çetin, et al. (ccetin2015depth). This approach directly determines the final position of each element by counting the number of elements less than it. Specifically, for an array of size $m$, the server performs m(m1)/2m(m-1)/2 homomorphic less-than operations and an equal number of homomorphic additions on the less-than operations’ result for computing the Hamming weight, which yields the correct indices $I$ for each element.
@@
where $S[i]$ is the element at the position $I$ of sorted array $S$ and $X$ is the input array. $\text{EQ}(i,I[j])$ is 1 if the $i$ equals the index $I[i]$ corresponding to $X[j]$, and 0 otherwise. Constructing one position $S[i]$ requires $m$ equality evaluation and $m$ multiplication, so sorting the entire encrypted array requires $O(m^{2})$ non-linear operations (less-than and equality evaluation) and $O(m^{2})$ linear operations (additions and multiplications).
Thus, homomorphically sorting an encrypted array of $m$ elements demands general FHE methods, particularly as the linear operations are performed on the results of non-linear operations, corresponding to our workload-2.
@@
Experimental Analysis. We conduct experiments using various general FHE methods on the private sorting tasks. Figure 7 (a) presents the performance of different FHE methods under different input bit precisions, and Figure 7 (b) shows the performance under different array dimensions. For instance, sorting an 8-element array requires 59.9 seconds with TFHE, compared to 645 seconds for Encoding Switching and 27,457 seconds for Scheme Switching. This efficiency advantage is due to two key factors. First, unlike other applications, private direct sorting on a single array cannot leverage SIMD (iliashenko2021faster), which diminishes the benefits of word-wise methods like Scheme Switching and Encoding Switching. Second, in this context, one of the operands in the linear operations is the output of a non-linear operation (either 0 or 1), resulting in a simpler circuit compared to linear operations on arbitrary integers. Conversely, Scheme Switching shows the lowest efficiency because its frequent switching, triggered by each non-linear operation following a linear one, adds significant overhead.
@@
The reasons are two-fold; firstly, unlike other applications, private direct sorting of an encrypted array cannot leverage the SIMD (iliashenko2021faster), which disables the advantages of word-wise Scheme Switching and Encoding Switching. Secondly, one of the linear operation operands is the result of a non-linear operation, i.e., 0 or 1. So that the TFHE circuit would be much less than linear operation between two random integers. On the other hand, Scheme Switching presents the lowest efficiency across all input bits and array sizes because of the frequent switching, since every non-linear operation will follow a linear operation.
@@ 7.6. Private Database Aggregation
Private Database Aggregation (PDBA) is a protocol that protects client privacy during aggregate queries on cloud-stored databases, such as Microsoft Azure SQL Server (mircroSQL) and AWS Aurora (AWSAurora). Recently, Leidos has partnered with AWS to explore the solution of using FHE to protect databases (fhe_database). It allows clients to securely store encrypted data on a cloud server and perform operations, like summation, without disclosing data or the details of the queries. The server processes these functions on encrypted data and returns only the encrypted aggregate results, ensuring the server does not access actual data or query specifics.
Recently, researchers have developed encrypted databases to ensure data privacy while enabling encrypted data processing. Cryptography-based solutions (popa2011cryptdb; tu2013processing; papadimitriou2016big; hackenjos2020sagma), utilize cryptographic primitives such as DET (bellare2007deterministic), SE (cash2013highly; curtmola2006searchable), partially PHE (paillier1999public; rivest1978method) and order OPE/ORE (chenette2016practical; boldyreva2009order). Recent advancements have explored the application of FHE in HEDA (ren2022heda), HE3DB (bian2023he3db), ArcEDB (zhang2024arcedb), and Engorgio (bian2025engorgio).
@@
In this query, the filtering predicates ”salary * work_hours BETWEEN 5000 AND 6000” and ”salary + bonus BETWEEN 700 AND 800” involve a non-linear operation following a linear operation. This requirement necessitates general FHE methods and aligns with workload-1 defined in Section 6.
While word-wise FHE methods work well for linear operations, they are less effective for non-linear operations like comparisons and logic filtering. To address this limitation and leverage the high efficiency of SIMD, Ren et al. introduce a novel framework, HEDA (ren2022heda). The core idea is to utilize scheme switching between RLWE and LWE ciphertexts. This approach conducts non-linear operations on LWE ciphertexts, where such operations are more naturally supported. Then, it converts the results back to RLWE ciphertexts to perform linear operations in a SIMD fashion.
Experimental Analysis. As described in Section 7.6, SQL queries involve combinations of non-linear and linear operations. To evaluate the performance of each general FHE method on the private database query, we conducted experiments on Query 1 across databases with different numbers of rows under 8-bit precision, as illustrated in Figure 8 (b).
The results show that the word-wise general FHE methods generally outperform the bit-wise TFHE, largely due to their use of SIMD capabilities, a feature that the bit-wise TFHE lacks. Furthermore, as the number of rows increases, Encoding Switching demonstrates better efficiency than Scheme Switching. For example, Encoding Switching takes 115 seconds for a 512-row database, compared to 1,626 seconds for Scheme Switching. This advantage is attributed to Encoding Switching’s ability to perform non-linear operations in a SIMD-style, while Scheme Switching processes non-linear operations on bit-wise ciphertexts and cannot leverage SIMD.
@@ 7.7. Private Neural Network
One of the most valuable applications of FHE is in facilitating privacy-preserving deep learning, i.e., Deep Learning as a Service (DLaaS) (srinivasan2019delphi; mohassel2017secureml), which is commonly used in information-sensitive areas such as financial (han2019logistic) and healthcare (kim2015private). This scenario involves a server that possesses a deep learning model, i.e., a neural network, and a client who encrypts their input data before sharing it with the untrusted server. The server processes encrypted input to provide predictions without accessing any sensitive information, thereby ensuring a secure inference service.
@@
Neural networks consist of linear layers such as convolutions and non-linear layers like ReLU and MaxPool (krizhevsky2012imagenet; lecun1998gradient). The combination of linear and non-linear operations required by these layers necessitates general FHE methods; indeed, the iterative sequence of linear and non-linear layers forms the basis of workload-3 as defined in Section 6.
@@
Despite the use of general FHE methods, some research leveraged the specific feature of neural networks to improve efficiency, but sacrificed some accuracy. Specifically, polynomial approximation-based methods (lee2022privacy; lee2023precise; lee2022optimization) leverage the noise resilience of neural networks to enhance efficiency while reducing multiplicative depth and complexity. Although these approximations are typically accurate only within a narrow range (often between -1 and 1), activation layer inputs are usually constrained (e.g., within [32,32][-32,32] or [64,64][-64,64]). By scaling these inputs to fit within [1,1][-1,1] using a predetermined scalar, servers can effectively implement these approximation techniques throughout the network, as demonstrated in approaches like DaCapo (cheondacapo), OPT-CNN (kim2023optimized), and Hetal (lee2023hetal), among others (lee2022low; dathathri2020eva). However, determining the appropriate scaling factor is critical: if it is too large, inputs may not be sufficiently constrained within [1,1][-1,1]; if too small, the inputs become overly diminished, resulting in significant approximation errors near 0.
@@
Another approximation strategy simulates non-linear activation functions using linear operations. For example, CryptoNets (gilad2016cryptonets) replaces ReLU with the square function, a method later adopted by frameworks like LoLa (brutzkus2019low) and Falcon (lou2020falcon). This approach avoids the overhead of homomorphic non-linear operations by training the neural network with a substitute activation function.
@@
Experimental Analysis. To evaluate the performance of various private neural network inference implementations described in Section 7.7, we conducted experiments using the CIFAR-10 dataset (krizhevsky2009learning) on a VGG-9 model. VGG-9 is a streamlined version of the VGG network family, comprising 9 layers including convolutional and fully connected layers. Typically, it consists of 6 convolutional layers followed by 3 fully connected layers, with ReLU activations and max pooling operations interleaved between the convolutional layers. Encryption parameters were set as specified in each referenced paper to ensure at least 128-bit security.
@@
Figure 8 (a) shows that implementations using the square function as an activation substitute, such as Falcon (lou2020falcon) and LoLa (brutzkus2019low), yield lower performance. This is primarily because training with degree-2 polynomial activations can lead to instability issues like exploding or vanishing gradients (ali2020polynomial; santriaji2024dataseal; xue2022audit; xue2025dictpflefficientprivatefederated; zheng2023cryptography). Although the square function enables faster inference, it is generally only suitable for smaller, less complex neural networks.
@@
In contrast, private neural networks that employ approximation-based activation functions, such as OPP-CNN (kim2023optimized) and EVA (dathathri2020eva), achieve higher accuracy levels. Meanwhile, TFHE’s performance is significantly slower compared to implementations using interpolation approximations. This discrepancy arises because neural network architectures predominantly consist of linear layers, which benefit from the SIMD capabilities of word-wise CKKS schemes. Bit-wise TFHE, lacking SIMD support and optimized linear operations, exhibits lower performance. Additionally, the Scheme Switching method (SHE (lou2019she)) achieves the highest accuracy but at slower inference times than the approximation-based approaches.
@@ 8. Recommendation and Discussion
First, if the application involves only linear operations, word-wise FHE methods (BGV, BFV, CKKS) are the most efficient choice. Conversely, if the application requires only non-linear operations, the next consideration should be whether the application can benefit from the word-wise SIMD technique. If SIMD does not offer an advantage, a bit-wise FHE method, such as TFHE, is the most efficient option. However, if SIMD is useful, several techniques enable non-linear functionalities in a word-wise context, including direct polynomial interpolation (narumanchi2017performance), decomposition-based polynomial interpolation (iliashenko2021faster), general functional bootstrapping (geelen2023bootstrapping), or special encoding methods like XCMP (akhavan2023level). These word-wise methods enable parallel computation to improve efficiency.
When an application demands both linear and non-linear computations (i.e., computational generality-required applications), the next question is whether the workload can benefit from SIMD parallelism. If SIMD does not offer advantages, e.g., private sorting for an array, a bit-wise FHE method, such as TFHE, remains the most practical choice. If SIMD is beneficial, the final consideration is whether exact, error-free results are necessary. If approximate results are acceptable, as is common in tasks like neural network inference, where noise tolerance is inherent, CKKS with approximation (cheon2020efficient; lee2021minimax) offers the most efficiency. If exact computation is required, the Encoding Switching (zhang2023hebridge) between beFV and FV, and the Scheme Switching (lu2021pegasus) between word-wise and bit-wise FHE can deliver the best performance.
While our SoK focuses on pure FHE methods for non-interactive general computation, we acknowledge related approaches with different trade-offs. Hybrid FHE-MPC protocols (xue2023cryptotrain; huang2022cheetah; he2024rhombus; hao2022iron; pang2024bolt) improve efficiency by using FHE for linear operations and MPC for non-linear operations. However, these hybrid approaches require frequent client-server communication during computation, sacrificing non-interactivity. In contrast, the pure FHE methods we focus on enable servers to complete the entire computation cycle independently, without requiring the client to remain online. This non-interactive property is critical for bandwidth-constrained environments.
Another research direction is optimizing rotation key management to reduce client-to-server transmission (cheon2025towards; lee2023rotation) or accelerate linear operations (zheng2023new; aikata2024secure) via specical encoding. While valuable, these works are orthogonal to our focus: we systematically analyze which FHE methods achieve functional generality supporting arbitrary combinations of linear and non-linear operations non-interactively, thereby establishing the fundamental computational capabilities and limitations of pure FHE for general AI workloads.